OpenStudy (anonymous):

use implicit differentiation to find dy/dx: xe^y-ye^x=1

OpenStudy (amistre64):

yay!!! implicits

OpenStudy (amistre64):

Dx(x e^y) = Dx(y e^x) = Dx(1)

OpenStudy (anonymous):

I LOVE implicits!

OpenStudy (amistre64):

x e^y y' + e^y - (y' e^x + e^x) = 0

OpenStudy (amistre64):

lol .... me too :)

OpenStudy (anonymous):

I am your fan

OpenStudy (amistre64):

aww shucks....ill fan you just so you can be cool like me :)

OpenStudy (amistre64):

x e^y y' + e^y - y' e^x - e^x = 0 if it aint got a y'...move it to the otherside

OpenStudy (amistre64):

y' (x e^y) - y'(e^x) = e^x - e^y

OpenStudy (amistre64):

factor out you y' and divide the rest...

OpenStudy (amistre64):

y' [(x e^y) - (e^x)] = e^x - e^y y' = e^x - e^y ---------- x e^y - e^x